OpenStudy (anonymous):

Solve for x: 3|2x - 2| + 6 = 18 x = 3, x = -3 x = 3, x = -1 x = 1, x = -1 x = -3, x = 1

OpenStudy (issy14):

3|2x-2|+6=18 6x-6+6=18 6x=18 x=18/6 x=3

OpenStudy (anonymous):

So A?

OpenStudy (issy14):

yup

OpenStudy (anonymous):

x=3